solve by elimination
4x + 4y = 5
2x - 8y = -5

Answer by mananth(16946)

4x + 4y = 5
2x - 8y = -5
4x+4y=5 .............1
2x-8y=-5 .............2
Eliminate y
multiply (1)by 4
Multiply (2) by 2
16x+16y = 20
4x-16y= -10
Add the two equations
20x=10
/20
x= 0.50
plug value of x in (1)
4x+4y= 5
2+4y=5
4y=5-2
4y=3
y=0.75